# Phase error rate estimation with basis-efficiency mismatch for decoy-state BB84

### v1

@@ IV Phase error estimation with basis-independent loss assumption
In this section, we will prove Eq. 7 for an implementation that satisfies the basis-independent loss assumption. It is useful to refer to Fig. 1 for this section. To prove Eq. 7, we will need to modify the actual protocol to an equivalent protocol (in the sense of being the same quantum to classical channel). To do so we will use Lemma 1 below to reformulate Alice and Bob’s measurements to consist of two steps. The first step will implement a basis-independent filtering operation that discards the inconclusive outcomes, while the second step will complete the measurement procedure. The the required claim will follow from random sampling arguments on the second step measurements. We start by explaining the two-step protocol measurements.
IV.1 Protocol Measurements
@@
We will first use the following lemma to divide Alice and Bob’s measurement procedure into two steps. For the proof, we refer the reader to Appendix A. We will use $S_{\bullet}(A)$ and $S_{\circ}(A)$ to denote the set of sub-normalized and noramlized states on the register $A$ respectively.
@@
Consider the POVMs $\{\Gamma_{(b_{A},b_{B}),(\neq)},\Gamma_{(b_{A},b_{B}),(=)},\Gamma_{(b_{A},b_{B}),(\bot)}\}$ defined in Eq. 3, which correspond to Bob obtaining a conclusive outcome (and Alice and Bob obtaining an error), Bob obtaining a conclusive outcome (and Alice and Bob not obtaining an error), and Bob obtaining an inconclusive outcome respectively, for basis choices $b_{A},b_{B}$. Without loss of generality, we can use Lemma 1 to equivalently describe Alice and Bob’s measurement procedure as consisting of two steps.
@@
IV.2 Constructing an equivalent protocol
@@
    This is then followed by random basis choices and assignment to test vs key by Alice and Bob. All $X$ basis rounds are used for testing, while some $Z$ basis rounds are used for key generation (and a small fraction is used to estimate $e^{\text{obs}}_{Z}$). Thus, we get the second node of Fig. 1. Note that the estimate $e^{\text{obs}}_{Z}$ of the error rate in the key bits is only used to determine the amount of error-correction required and does not affect the secrecy of the protocol. However, $e^{\text{obs}}_{Z}$ and the choice of error-correction protocol is important to ensure that error-verification succeeds with high probability.
@@
IV.3 Sampling
We will now turn our attention to the sampling part of the argument, and obtain an estimate $\mathcal{B}_{0,0}$ on the phase error rate that satisfies Eq. 7. To do so, we will make use of the following Lemma, which uses the Serfling bound [36]. For the proof, we refer the reader to Section B.1.
@@ IV.3 Sampling
There are two important aspects to the sampling argument. First, the Serfling bound applies in the situation where one chooses a random subset of fixed-length for testing. However, the above procedure (and many QKD protocols) randomly assigns each round to testing vs key generation. Thus, Serfling must be applied with some care, and that is what is done here (see footnote.  [^5] It is also worthwhile to note that if one is interested in estimating the QBER independent of basis, then the standard serfling argument is directly applicable (for instance in [2]).). This observation has been missing in many prior works. Second, since we are interested in a variable-length protocol, we require slightly different statements than standard fixed-length security proofs (Eq. 7). However, these can also be obtained by simple (almost trivial) modifications to existing arguments and yield the same results as before. Both these issues are addressed in the proof of Lemma 2 in Appendix B.
Let us consider the second node in the equivalent protocol constructed in Fig. 1, where rounds are now randomly assigned for testing ($X$ basis) or key generation ($Z$ basis and key generation). (The remaining rounds are used for estimating the $Z$ basis error rate or discarded and are unimportant for this discussion). Consider the state $\rho_{|\Omega_{(n_{X},n_{K})}}$, where the number of rounds to be used to testing and key generation is fixed. Using Lemma 2 from Appendix B on this state, we obtain
@@
Thus, for the above choice of $\mathcal{B}_{0,0}(e^{\text{obs}}_{X},n_{X},n_{K})$, the variable-length security of the protocol follows from Theorem 1.
IV.4 Comparing to fixed-length protocols
V Phase error estimation without basis-independent loss assumption
@@
Figure 2: Protocol flowchart for the equivalent protocol from Section V, where basis-independent loss assumption (Eq. 4) is not satisfied. The dotted arrows and boxes represent virtual measurements that do not actually happen in the real protocol. Connections between the error rates in different boxes are highlighted using curved arrows. We use $\bm{N}_{P}$ to denote the number of $P$ measurement outcomes, where $P$ denotes a POVM element. For the POVMs, the reader may refer to Table 1 or Section V.1. For clarity, we have omitted the conditioning on events in the figure (but not in our proof). Compared to Fig. 2, the testing and key generation rounds go through an additional second step filtering measurement that depends on the basis used, which typically results in a few rounds being discarded, before undergoing the final measurement. The basis used in these measurements in indicated in each box, and indicates the basis used by both Alice and Bob.
In this section, we will prove Eq. 7 for an implementation that does not satisfy the basis-independent loss assumption. The argument is similar to the one presented in Section IV, with important additions. It is helpful to refer to Fig. 2 for this section. We will first explain the idea behind the proof, before stating the proof itself.
@@
We will use Lemma 1 in Section V.1 to construct an equivalent measurement procedure (in the sense that it is the same quantum to classical channel) for the protocol, which consists of three steps. The first step measurement is done using the POVM $\{\tilde{F},\mathrm{I}-\tilde{F}\}$ and implements basis-independent filtering (discarding) operations. ($\tilde{F}$ here plays the same role as in Section IV, but is defined differently). In particular it is the largest common filtering operation over both basis choices.
@@
In particular, we will relate the error rates before and after discarding for the testing rounds ($\tilde{e}^{\text{obs}}_{XX}\leftrightarrow e^{\text{obs}}_{X}$) by simply noting that we discard rounds in the second-step measurements. On the other hand, we will relate $\tilde{e}^{\text{key}}_{ZX}\leftrightarrow e^{\text{key}}_{X}$ by bounding the number of discards that can happen in the second step filtering measurements. This relation will depend on $\delta_{2}$, which will be the metric that quantifies the “smallness” of the POVM element corresponding to the discard outcome. $\tilde{e}^{\text{obs}}_{XX}$ and $\tilde{e}^{\text{key}}_{XX}$ correspond to error rates corresponding to exactly the same measurement, and assigned to test vs key randomly. Thus, they can be related using Serfling (Lemma 2), exactly as in Section IV. $\tilde{e}^{\text{key}}_{XX}$ and $\tilde{e}^{\text{key}}_{ZX}$ correspond to error rates on the same state, but with slightly different POVMs, and thus are expected to be similar. This can be rigorously argued using Lemma 3, where we use $\delta_{1}$ to quantify the “closeness” of these POVMs. Combining all these relations, we will ultimately obtain Eq. 44.
@@
V.1 Protocol Measurements
Fix the basis $b_{A},b_{B}$ used by Alice and Bob. As in Section IV.1, consider the POVM $\{\Gamma_{(b_{A},b_{B}),(\neq)},\Gamma_{(b_{A},b_{B}),(=)},\Gamma_{(b_{A},b_{B}),(\bot)}\}$ defined in Eq. 3, which correspond to Bob obtaining a conclusive outcome (and Alice and Bob obtaining an error), Bob obtaining a conclusive outcome (and Alice and Bob not obtaining an error), and Bob obtaining an inconclusive outcome respectively. As in Section IV.1, we can use Lemma 1 to equivalently describe Alice and Bob’s measurement procedure as consisting of two steps. First, they measure using $\{\tilde{F}_{(b_{A},b_{B}),(\text{con})},\tilde{F}_{(b_{A},b_{B}),(\bot)}\}$ which determines whether they obtain a conclusive or inconclusive measurement outcome. Then, if they obtain a conclusive outcome, they measure using the POVM $\{G^{\text{con}}_{(b_{A},b_{B}),(=)},G^{\text{con}}_{(b_{A},b_{B}),(\neq)}\}$.
However, due to basis-efficiency mismatch, $\tilde{F}_{(b_{A},b_{B}),(\text{con})}$ will depend on the basis choices. Therefore we cannot proceed further in the same way as in Section IV. Instead, we choose an $\tilde{F}$ such that
$\tilde{F}_{(b_{A},b_{B}),(\text{con})}\leq\tilde{F}\quad\forall(b_{A},b_{B})$ (26)
@@ V.1 Protocol Measurements
Since basis-mismatch rounds are discarded anyway, it is possible to argue that we only need $\tilde{F}_{(b_{A},b_{B}),(\text{con})}\leq\tilde{F}$ for $b_{A}=b_{B}$. This involves constructing a slightly different equivalent protocol where the first node decides basis match vs mismatch. The basis match events then undergo the usual filtering followed by basis choice, while the mismatch events are discarded without any filtering. If this modified requirement results in a value of $\tilde{F}$ that is “smaller” then the original choice, then this will lead to tighter key rates. Intuitively, this is due to the fact that a smaller value of $\tilde{F}$ means that more loss is attributed to the basis-independent filtering.
We will now reformulate the first step measurement to include this basis-independent filtering. Consider the three outcome POVM given by $\{\mathrm{I}-\tilde{F},\tilde{F}-\tilde{F}_{(b_{A},b_{B}),(\text{con})},\tilde{F}_{(b_{A},b_{B}),(\text{con})}\}$. The first outcome corresponds to the basis-independent filter. Then, using Lemma 1 again, we can view this as two-step measurement. In the first step, Alice and Bob measure using $\{\tilde{F},\mathrm{I}-\tilde{F}\}$. If they obtains the $\tilde{F}$ outcome, they measure a second time using $\{{F}_{(b_{A},b_{B}),(\text{con})},\mathrm{I}-{F}_{(b_{A},b_{B}),(\text{con})}\}$ which does the basis-dependent part of the filtering. Thus for any choice of $(b_{A},b_{B})$, there are three different measurement steps in the protocol. Their exact procedure is clarified below in the construction of the equivalent protocol from Fig. 2. We also provide Table 1 which includes a brief description of the POVMs.
@@
It will be convenient to combine the second and third step measurement into a single measurement step with three outcomes. For brevity we introduce the following notation to write this POVM $\{G^{\text{con,F}}_{(b_{1}b_{2}),(\neq)},G^{\text{con,F}}_{(b_{1}b_{2}),(=)},\mathrm{I}-{F}_{(b_{1}),(\text{con})}\}$ where
@@
where the subscript $b_{1}b_{2}$ determines the basis for the second step and third step measurements by both Alice and Bob, and the superscript $F$ indicates the merging of the two measurement steps.
@@
V.2 Constructing an equivalent protocol
We will now construct the equivalent protocol from Fig. 2. The construction is similar to the one from Section IV.2, albeit with some important modifications.
@@ V.2 Constructing an equivalent protocol
    As in Section IV.2, we observe that the first step measurement is conducted using $\{\tilde{F},\mathrm{I}-\tilde{F}\}$ and is independent of basis. Therefore, we can delay basis choice until after this measurement has been completed, and the $\mathrm{I}-\tilde{F}$ outcomes are discarded. That is the first node of Fig. 2.
@@
    The remaining rounds undergo random basis choice. Basis mismatch rounds are discarded, all $X$ basis rounds are used for testing, while $Z$ basis rounds are probabilistically chosen for testing and key generation. This allows us to obtain the second node of Fig. 2. Again, as in Section IV.2, the estimate we obtain on $e^{\text{obs}}_{Z}$ does not affect the secrecy claim of the protocol, since $e^{\text{obs}}_{Z}$ is only used to determine the amount of error-correction to be performed.
Note that unlike Section IV.2, we have to perform two measurements on the testing and key generation rounds after the second node, and these rounds are not guaranteed to result in a conclusive outcome. We describe these measurements in detail below.
V.2.1 Testing Rounds after basis-independent Filter

### v2

@@ III Phase error estimation with basis-independent loss assumption
In this section, we will prove Eq. 5 for an implementation that satisfies the basis-independent loss assumption. It is useful to refer to Fig. 2 for this section. To prove Eq. 5, we will need to modify the actual protocol to an equivalent protocol (in the sense of being the same quantum to classical channel). To do so we will use Lemma 1 below to reformulate Alice and Bob’s measurements to consist of two steps. The first step will implement a basis-independent filtering operation that discards the inconclusive outcomes, while the second step will complete the measurement procedure. Then the required claim will follow from random sampling arguments on the second step measurements. We start by explaining the two-step protocol measurements.
III.1 Protocol Measurements
@@
We will first use the following lemma to divide Alice and Bob’s measurement procedure into two steps. For the proof, we refer the reader to Appendix A. We will use $S_{\bullet}(A)$ and $S_{\circ}(A)$ to denote the set of sub-normalized and normalized states on the register $A$ respectively.
@@
$\{\Gamma_{k}\}_{k\in\mathcal{A}}$
$\{G_{k}\}_{k\in\mathcal{A}_{i}}$
[…]
$\Leftrightarrow$
Figure 1: Schematic for the two-step measurement procedure from Lemma 1. Note that the second step measurement $\{G_{k}\}_{k\in\mathcal{A}_{i}}$ depends on the outcome of the first step measurement.
Consider the POVMs $\{\Gamma_{(b_{A},b_{B}),(\neq)},\Gamma_{(b_{A},b_{B}),(=)},\Gamma_{(b_{A},b_{B}),(\bot)}\}$ defined in Eq. 3, which correspond to Bob obtaining a conclusive outcome and Alice and Bob obtaining an error, Bob obtaining a conclusive outcome and Alice and Bob not obtaining an error, and Bob obtaining an inconclusive outcome respectively, for basis choices $b_{A},b_{B}$. Without loss of generality, we can use Lemma 1 to equivalently describe Alice and Bob’s measurement procedure as consisting of two steps.
@@
III.2 Constructing an equivalent protocol
[…]
Figure 2: Protocol flowchart for the equivalent protocol from Section III, where basis-independent loss assumption (Eq. 4) is satisfied. The dotted arrows and boxes represent virtual measurements that do not actually happen in the real protocol. Connections between different boxes are highlighted using curved arrows. We use the Serfling bound (Lemma 2) to obtain a bound on the phase error rate from observations. The phase error rate is then used to bound the smooth min entropy using the EUR statement. We use $\bm{N}_{P}$ to denote the number of $P$ measurement outcomes, where $P$ denotes a POVM element. For clarity, we have omitted the conditioning on events in the figure (but not in our proof). The basis used for measurements is indicated in each box, and refers to the basis used by both Alice and Bob.
@@ III.2 Constructing an equivalent protocol
    This is then followed by random basis choices and assignment to test vs key by Alice and Bob. All $X$ basis rounds are used for testing, while most $Z$ basis rounds are used for key generation (and a small fraction is used to estimate $e^{\text{obs}}_{Z}$). Thus, we get the second node of Fig. 2. Note that the estimate $e^{\text{obs}}_{Z}$ of the error rate in the key bits is only used to determine the amount of error-correction required and does not affect the secrecy of the protocol. However, $e^{\text{obs}}_{Z}$ and the choice of error-correction protocol is important to ensure that error-verification succeeds with high probability.
@@
III.3 Sampling
We will now turn our attention to the sampling part of the argument, and obtain an estimate $\mathcal{B}_{0,0}$ on the phase error rate that satisfies Eq. 5. To do so, we will make use of the following Lemma, which uses the Serfling bound [36]. For the proof, we refer the reader to Section C.1.
@@ III.3 Sampling
There are two important aspects to the sampling argument. First, the Serfling bound applies in the situation where one chooses a random subset of fixed-length for testing. However, the above procedure (and many QKD protocols) randomly assigns each round to testing vs key generation. Thus, Serfling must be applied with some care, and that is what is done here (see footnote.  [^4] It is also worthwhile to note that if one is interested in estimating the QBER independent of basis, then the standard serfling argument is directly applicable (for instance in [2]).). This observation has been missing in many prior works. Second, since we are interested in a variable-length protocol, we require slightly different statements than standard fixed-length security proofs (Eq. 5). However, these can also be obtained by simple (almost trivial) modifications to existing arguments and yield the same results as before. Both these issues are addressed in the proof of Lemma 2 in Appendix C.
Let us consider the second node in the equivalent protocol constructed in Fig. 2, where rounds are now randomly assigned for testing ($X$ basis) or key generation ($Z$ basis and key generation). (The remaining rounds are used for estimating the $Z$ basis error rate or discarded and are unimportant for this discussion). Consider the state $\rho_{|\Omega_{(n_{X},n_{K})}}$, where the number of rounds to be used to testing and key generation is fixed. Using Lemma 2 on this state, we obtain
@@
(In this work, we will use the convention that $\sum_{x}$ denotes the sum over all possible values $x$ can take). Thus, for the above choice of $\mathcal{B}_{0,0}(e^{\text{obs}}_{X},n_{X},n_{K})$, the variable-length security of the protocol follows from Theorem 1.
IV Phase error estimation without basis-independent loss assumption
@@
Figure 3: Protocol flowchart for the equivalent protocol from Section IV, where basis-independent loss assumption (Eq. 4) is not satisfied. The dotted arrows and boxes represent virtual measurements that do not actually happen in the real protocol. Connections between the error rates in different boxes are highlighted using curved arrows. We use $\bm{N}_{P}$ to denote the number of $P$ measurement outcomes, where $P$ denotes a POVM element. For the POVMs, the reader may refer to Table 1 or Section IV.1. For clarity, we have omitted the conditioning on events in the figure (but not in our proof). Compared to Fig. 3, the testing and key generation rounds go through an additional second step filtering measurement that depends on the basis used, which typically results in a few rounds being discarded, before undergoing the final measurement. The basis used in these measurements in indicated in each box, and indicates the basis used by both Alice and Bob.
In this section, we will prove Eq. 5 for an implementation that does not satisfy the basis-independent loss assumption. The argument is similar to the one presented in Section III, with important additions. It is helpful to refer to Fig. 3 for this section. We will first explain the idea behind the proof, before stating the proof itself.
@@
We will use Lemma 1 in Section IV.1 to construct an equivalent measurement procedure (in the sense that it is the same quantum to classical channel) for the protocol, which consists of three steps. The first step measurement is done using the POVM $\{\tilde{F},\mathrm{I}-\tilde{F}\}$ and implements basis-independent filtering (discarding) operations. ($\tilde{F}$ here plays the same role as in Section III, but is defined differently). In particular it is the largest common filtering operation over both basis choices.
@@
In particular, we will relate the error rates before and after discarding for the testing rounds ($\tilde{e}^{\text{obs}}_{XX}\leftrightarrow e^{\text{obs}}_{X}$) by simply noting that we discard rounds in the second-step measurements. On the other hand, we will relate $\tilde{e}^{\text{key}}_{ZX}\leftrightarrow e^{\text{key}}_{X}$ by bounding the number of discards that can happen in the second step filtering measurements. This relation will depend on $\delta_{2}$, which will be the metric that quantifies the “smallness” of the POVM element corresponding to the discard outcome. $\tilde{e}^{\text{obs}}_{XX}$ and $\tilde{e}^{\text{key}}_{XX}$ correspond to error rates corresponding to exactly the same measurement, and assigned to test vs key randomly. Thus, they can be related using Serfling (Lemma 2), exactly as in Section III. $\tilde{e}^{\text{key}}_{XX}$ and $\tilde{e}^{\text{key}}_{ZX}$ correspond to error rates on the same state, but with slightly different POVMs, and thus are expected to be similar. This can be rigorously argued using Lemma 3, where we use $\delta_{1}$ to quantify the “closeness” of these POVMs. Combining all these relations, we will ultimately obtain Eq. 34.
@@
IV.1 Protocol Measurements
Fix the basis $b_{A},b_{B}$ used by Alice and Bob. As in Section III.1, consider the POVM $\{\Gamma_{(b_{A},b_{B}),(\neq)},\Gamma_{(b_{A},b_{B}),(=)},\Gamma_{(b_{A},b_{B}),(\bot)}\}$ defined in Eq. 3, which correspond to Bob obtaining a conclusive outcome (and Alice and Bob obtaining an error), Bob obtaining a conclusive outcome (and Alice and Bob not obtaining an error), and Bob obtaining an inconclusive outcome respectively. Since $\Gamma_{(b_{A},b_{B}),(\bot)}$ now depends on the basis choices, we cannot proceed in the same way as before. This reflects the fact that the discarding is basis dependent. Thus we will reformulate the measurement process in a different way.
To do so, consider a $\tilde{F}$ such that
$\tilde{F}\geq\Gamma_{(b_{A},b_{B}),(=)}+\Gamma_{(b_{A},b_{B}),(\neq)}\quad\forall(b_{A},b_{B})$ (16)
@@ IV.1 Protocol Measurements
Since basis-mismatch rounds are discarded anyway, it is possible to argue that we only need $\tilde{F}$ to satisfy $\Gamma_{(b_{A},b_{B}),(=)}+\Gamma_{(b_{A},b_{B}),(\neq)}\leq\tilde{F}$ for $b_{A}=b_{B}$. This involves constructing a slightly different equivalent protocol where the first node decides basis match vs mismatch. The basis match events then undergo the usual filtering followed by basis choice, while the mismatch events are discarded without any filtering. If this modified requirement results in a value of $\tilde{F}$ that is “smaller” then the original choice, then this will lead to tighter key rates. Intuitively, this is due to the fact that a smaller value of $\tilde{F}$ means that more loss is attributed to the basis-independent filtering.
To reformulate the measurement procedure, start by considering the four-outcome POVM given by $\{\mathrm{I}-\tilde{F},\tilde{F}-\Gamma_{(b_{A},b_{B}),(=)}-\Gamma_{(b_{A},b_{B}),(\neq)},\Gamma_{(b_{A},b_{B}),(=)},\Gamma_{(b_{A},b_{B}),(\neq)}\}$, where the first two outcomes correspond to discard, the third correspond to a conclusive no-error outcome, and the fourth corresponds to a conclusive error. This four-outcome measurement followed by classical grouping of the first two outcomes is then equivalent to the original three-outcome measurement in the protocol.
Now, we can use Lemma 1 to reformulate the four-outcome measurement as occurring in two steps. In the first step, Alice and Bob measure using POVM $\{\tilde{F},\mathrm{I}-\tilde{F}\}$ and discard the latter outcomes. If they obtain the $\tilde{F}$ outcome, they then complete the measurement using POVM $\{\tilde{F}_{(b_{A},b_{B}),(\bot)},\tilde{F}_{(b_{A},b_{B}),(=)},\tilde{F}_{(b_{A},b_{B}),(\neq)}\}$, corresponding to discard, conclusive no-error and conclusive error outcomes respectively.
We then use Lemma 1 again to reformulate this three-outcome measurement to consist of two steps. First, they measure using $\{{F}_{(b_{A},b_{B}),(\text{con})},{F}_{(b_{A},b_{B}),(\bot)}\}$ which determines whether they obtain a conclusive or inconclusive measurement outcome. Then, if they obtain a conclusive outcome, they measure using the POVM $\{G^{\text{con}}_{(b_{A},b_{B}),(=)},G^{\text{con}}_{(b_{A},b_{B}),(\neq)}\}$. Thus we now have a three-step measurement procedure, described in Table 1.
@@
It will be convenient to recombine the second and third step measurement into a single measurement step with three outcomes. For brevity we introduce the following notation to write this POVM $\{G^{\text{con,F}}_{(b_{1}b_{2}),(\neq)},G^{\text{con,F}}_{(b_{1}b_{2}),(=)},\mathrm{I}-{F}_{(b_{1}),(\text{con})}\}$ where
@@
where the subscript $b_{1}b_{2}$ determines the basis for the second step and third step measurements by both Alice and Bob, and the superscript $F$ indicates the merging of the two measurement steps. (Note that if $b_{1}=b_{2}=b$, then this simply reverses the earlier action of Lemma 1 that split $\{\tilde{F}_{(b,b),(\bot)},\tilde{F}_{(b,b),(=)},\tilde{F}_{(b,b),(\neq)}\}$ to generate the second and third-step measurements. However, we will consider fictitious measurements where $b_{1}\neq b_{2}$ in our proof. To describe such measurements, it is indeed necessary to split $\{\tilde{F}_{(b_{A},b_{B}),(\bot)},\tilde{F}_{(b_{A},b_{B}),(=)},\tilde{F}_{(b_{A},b_{B}),(\neq)}\}$ into two separate steps.)
@@
IV.2 Constructing an equivalent protocol
We will now construct the equivalent protocol from Fig. 3. The construction is similar to the one from Section III.2, albeit with some important modifications.
@@ IV.2 Constructing an equivalent protocol
    As in Section III.2, we observe that the first step measurement is conducted using $\{\tilde{F},\mathrm{I}-\tilde{F}\}$ and is independent of basis. Therefore, we can delay basis choice until after this measurement has been completed, and the $\mathrm{I}-\tilde{F}$ outcomes are discarded. That is the first node of Fig. 3.
@@
    The remaining rounds undergo random basis choice. Basis mismatch rounds are discarded, all $X$ basis rounds are used for testing, while $Z$ basis rounds are probabilistically chosen for testing and key generation. This allows us to obtain the second node of Fig. 3. Again, as in Section III.2, the estimate we obtain on $e^{\text{obs}}_{Z}$ does not affect the secrecy claim of the protocol, since $e^{\text{obs}}_{Z}$ is only used to determine the amount of error-correction to be performed.
Note that unlike Section III.2, we have to perform two measurements on the testing and key generation rounds after the second node, and these rounds are not guaranteed to result in a conclusive outcome. We describe these measurements in detail below.
IV.2.1 Testing Rounds after basis-independent Filter